# A Practical Investigation of Spatially-Controlled Image Generation with Transformers

### v1

@@ 6.2 Results
Based on the exploration of the previous section, we select a set of sampling parameters that aim to balance generation quality with control consistency for our remaining experiments. They are specified in Tabs. 1 and 7. Additional generation examples for the following experiments can be found in Appendix B.
model method ctrl
@@
ctrl-G =1.0 prefill adapter depth 310M + 202M 5.64 208 29.7 0.31 12
@@
VAR-d16 CFG=3.0, temp=1, top-$p$=0.6, top-$k$=900 ctrl-G =1.5 prefill adapter depth 310M + 202M 5.47 201 27.9 0.56 8.5
model method ctrl
@@
ctrl-G =1.0 prefill adapter depth 675M + 446M 4.62 251 32.7 3.4 0.35
@@
SiT-XL/2 CFG=3.0, Euler -ODE, steps=64, proj-G ctrl-G =1.5 prefill adapter depth 675M + 446M 5.17 227 29.7 5.8 0.24
Table 1: Performance of adapter-based approaches compared to the prefill baseline. 1) Adapters have worse control consistency (F1,RMSE) compared to prefill+finetune, although generation quality is comparable. 2) The addition adapter is better for canny conditioning, whilst the prefill adapter is better for depth conditioning. 3) Adjusting sampling parameters (such as ctrl-G) can make up some of the difference in control consistency. 4) The prefill adapter is a little more parameter efficient (no zero layers) and considerably more inference efficient than the addition adapter. 5) Adapters enjoy no parameter overhead for ctrl-G, as the control-free base model weights are frozen and available.
@@
Tab. 1 shows results comparing the performance of the two adapter approaches to the prefill + finetuning baseline. We find that generally speaking adapters have worse control consistency, but have comparable generation quality. Intuitively, an adapter is less flexible to influence the generation process, limiting its ability to closely follow conditioning, but also retaining the original model’s generation capabilities. This aligns with existing research on LoRA in language modelling (Biderman et al., 2024). The addition adapter has the advantage for canny conditioning, whilst the prefill adapter is superior for depth conditioning, although both are worse than full finetuning. We also demonstrate that adjusting the sampling parameters of an adapter (in this case ctrl-G) can make up some of difference in control consistency, although the associated tradeoffs discussed in Sec. 5 still exist. We note that this suggests the comparisons made in ControlAR (Li et al., 2025) and Omnigen (Xiao et al., 2025) may not be entirely fair, as their approaches train all parameters, but they compare against adapter-based approaches.
@@
We find that the prefill adapter is slightly more parameter efficient (due to not needing any additional zero-initialised layers), although it is of course possible to explore other ways of improving parameter efficiency, such as reducing the depth/number of blocks in the adapter or using LoRA (Hu et al., 2022). We note that prefilling (adapter or not) is considerably more inference efficient compared to the addition adapter based on ControlNet (Zhang et al., 2023). The former is able to utilise a KV-cache, meaning that inference over control tokens only happens once per generation, whilst in the latter case, each generation token forward pass is paired with a spatially aligned control token forward pass, significantly increasing inference costs. This difference is especially apparent for SiT, where the KV-cache can be re-used for all (64) denoising steps. We note that Tan et al. (2025a; b) make a similar observation when designing their method for conditional generation built around FLUX (Labs, 2024).
@@
We also note that since adapters keep the original (no control) model’s weights intact, there is no parameter overhead for ctrl-G. As such it can be readily used and considered for many existing adapter-based systems.
@@
In order to investigate the phenomenon of “forgetting” we finetune and train adapters on a semantically pruned version of ImageNet where all samples with labels that are children of the WordNet (Miller et al., 1990) node “organism” are removed. We train for the same number of iterations as previously ($\sim$ 50K). Results are shown in Figs. 7 and 9. We find that training all VAR parameters exhibits a collapse in generation quality, where the model “forgets” how to generate “organism” classes outside of the control finetuning data. Conversely, we observe that adapters are able to mitigate this effect, successfully generalising control-conditioned generation to classes unseen in the control data (although present in the base model’s pretraining data). SiT interestingly is able to maintain its generation quality compared to finetuning on the full dataset. We hypothesise this robustness may be due to the larger size (675M vs 310M parameters) and longer pretraining (800 vs 200) of the SiT model (Yu et al., 2025) compared to the VAR model (Tian et al., 2024).
@@ A.1 Distribution Truncation
Fig. 8 shows how distribution truncation and CFG interact when ctrl-G=1.5. It tells a similar story to Fig. 4 and also illustrates how both distribution truncation and control guidance can be adjusted together at inference time to improve control consistency. For VAR, the control consistency improvements from ctrl-G and top-$p$ stack together.
@@
A.2 Learning with Limited Data
@@ Appendix B Example Generations
We include additional example generations over various experimental settings in Figs. 10, 11, 12 and 13. Readers can observe the effect of ctrl-G (canny: facial features of the dog, depth: seeds/stones next to the bird) across Figs. 10 and 12. Fig. 11 show further examples of VAR “forgetting” “organism” concepts from ImageNet-pretraining. We note these are more subtle for depth (plumage on bird, stripes rather than spots on salamander).
@@
Figure 11: Additional examples of VAR generations with limited control data – “organism” classes excluded. (CFG=3.0, temp=1, top-$p$=0.6, top-$k$=900, ctrl-G=1.0)

### v2

@@ 6.2 Results
Based on the exploration of the previous section, we select a set of sampling parameters that aim to balance generation quality with control consistency for our remaining experiments. They are specified in Tabs.˜1 and 7. Additional generation examples for the following experiments can be found in Appendix˜B.
model
(sampling params)
method ctrl
@@
ctrl-G
=1.0
prefill adapter depth 310M + 202M 5.64 208 29.7 0.31 12
@@
VAR-d16
CFG=3.0,
temp=1,
top-pp=0.6,
top-kk=900
ctrl-G
=1.5
prefill adapter depth 310M + 202M 5.47 201 27.9 0.56 8.5
(a)(b)(c)(d)(e)
model
(sampling params)
method ctrl
@@
ctrl-G
=1.0
prefill adapter depth 675M + 446M 4.62 251 32.7 3.4 0.35
@@
SiT-XL/2
CFG=3.0,
Euler
-ODE,
steps=64,
proj-G
ctrl-G
=1.5
prefill adapter depth 675M + 446M 5.17 227 29.7 5.8 0.24
(a)(b)(c)(d)(e)
Table 1: Performance of adapter-based approaches compared to the prefill baseline. (a) Adapters have worse control consistency (F1,RMSE) compared to prefill+finetune, although generation quality is comparable. (b) The addition adapter is better for canny conditioning, whilst the prefill adapter is better for depth conditioning. (c) Adjusting sampling parameters (such as ctrl-G) can make up some of the difference in control consistency. (d) The prefill adapter is a little more parameter efficient (no zero layers) and considerably more inference efficient than the addition adapter. (e) Adapters enjoy no parameter overhead for ctrl-G, as the control-free base model weights are frozen and available.
@@
Tab.˜1 shows results comparing the performance of the two adapter approaches to the prefill + finetuning baseline. We find that generally speaking adapters have worse control consistency, but have comparable generation quality. Intuitively, an adapter is less flexible to influence the generation process, limiting its ability to closely follow conditioning, but also retaining the original model’s generation capabilities. This aligns with existing research on LoRA in language modelling (Biderman et al., 2024). The addition adapter has the advantage for canny conditioning, whilst the prefill adapter is superior for depth conditioning, although both are worse than full finetuning. We also demonstrate that adjusting the sampling parameters of an adapter (in this case ctrl-G) can make up some of difference in control consistency, although the associated tradeoffs discussed in Sec.˜5 still exist. We note that this suggests the comparisons made in ControlAR (Li et al., 2025) and Omnigen (Xiao et al., 2025) may not be entirely fair, as their approaches train all parameters, but they compare against adapter-based approaches.
@@
We find that the prefill adapter is slightly more parameter efficient (due to not needing any additional zero-initialised layers), although it is of course possible to explore other ways of improving parameter efficiency, such as reducing the depth/number of blocks in the adapter or using LoRA (Hu et al., 2022). [^7] Sec. A.2 contains a supplementary experiment that shows that LoRA can be used to improve parameter efficiency at the cost of performance. We note that prefilling (adapter or not) is considerably more inference efficient compared to the addition adapter based on ControlNet (Zhang et al., 2023). The former is able to utilise a KV-cache, meaning that inference over control tokens only happens once per generation, whilst in the latter case, each generation token forward pass is paired with a spatially aligned control token forward pass, significantly increasing inference costs. This difference is especially apparent for SiT, where the KV-cache can be re-used for all (64) denoising steps. We note that Tan et al. (2025a; b) make a similar observation when designing their method for conditional generation built around FLUX (Labs, 2024).
@@
We also note that since adapters keep the original (no control) model’s weights intact, there is no parameter overhead for ctrl-G. As such it can be readily used and considered for many existing adapter-based systems. In Appendix˜C we demonstrate this, showing that applying ctrl-G to ControlNet++ (Li et al., 2024a) out of the box results in similar qualitative and quantitative trade offs to Sec.˜5.1 without any need for additional training (Fig.˜14).
@@
In order to investigate the phenomenon of “forgetting” we finetune and train adapters on a semantically pruned version of ImageNet where all samples with labels that are children of the WordNet (Miller et al., 1990) node “organism” are removed. We train for the same number of iterations as previously ($\sim$ 50K). Results are shown in Figs.˜7 and 9. We find that training all VAR parameters exhibits a collapse in generation quality, where the model “forgets” how to generate “organism” classes outside of the control finetuning data. Conversely, we observe that adapters are able to mitigate this effect, successfully generalising control-conditioned generation to classes unseen in the control data (although present in the base model’s pretraining data). SiT interestingly is able to maintain its generation quality compared to finetuning on the full dataset. We hypothesise this robustness may be due to the larger size (675M vs 310M parameters) and longer pretraining (800 vs 200) of the SiT model (Yu et al., 2025) compared to the VAR model (Tian et al., 2024).
@@ A.1 Distribution Truncation
Fig.˜8 shows how distribution truncation and CFG interact when ctrl-G=1.5. It tells a similar story to Fig.˜4 and also illustrates how both distribution truncation and control guidance can be adjusted together at inference time to improve control consistency. For VAR, the control consistency improvements from ctrl-G and top-$p$ stack together.
@@
A.2 Low Rank Adaptation (LoRA)
model
(sampling params)
method ctrl
#params
base + adapter
metrics
[…]
VAR-d16
CFG=3.0, temp=1,
top-pp=0.6, top-kk=900,
ctrl-G=1.0
prefill canny 310M 5.12 187 36.9
prefill + LoRA (r=64) canny 310M + 10M 5.57 185 33.2
prefill adapter canny 310M + 202M 4.98 201 33.0
prefill adapter + LoRA (r=64) canny 310M + 10M 5.59 182 28.9
Table 2: Experiments with Low-Rank Adaptation (LoRA). We use rsLoRA (Kalajdzievski, 2023) and set $r=\alpha=64$. We attach LoRA adapters to the QKVQKV matrices in each attention layer as well as the weight matrices in the feed-forward network. For prefill + LoRA the trainable weights see all tokens, whilst for prefill adapter + LoRA the trainable weights only see control tokens (Fig.˜6 where the trainable parameters are replaced with LoRA). LoRA enables much lower parameter overhead, but the reduced model flexibility comes at the cost of generation quality and control consistency.
Tab.˜2 shows an additional ablation, replacing the trainable parameters in Fig.˜6 with LoRA (Hu et al., 2022) adapters. We find that LoRA can considerably improve the parameter efficiency of learning a spatial control; however, it comes at the cost of generation quality as well as control consistency. This aligns with the behaviour reported in Li et al. (2025).
A.3 Learning with Limited Data
@@ Appendix B Example Generations
We include additional example generations over various experimental settings in Figs.˜10, 11, 12 and 13. Readers can observe the effect of ctrl-G (canny: facial features of the dog, depth: seeds/stones next to the bird) across Figs.˜10 and 12. Fig.˜11 show further examples of VAR “forgetting” “organism” concepts from ImageNet-pretraining. We note these are more subtle for depth (plumage on bird, stripes rather than spots on salamander).
@@
Figure 11: Additional examples of VAR generations with limited control data – “organism” classes excluded. (CFG=3.0, temp=1, top-pp=0.6, top-kk=900, ctrl-G=1.0)
@@
Appendix C Ctrl-G on Pre-trained Adapters for T2I Models
We demonstrate that ctrl-G can be directly applied to existing opensource adapters without the need for any additional training. We show the effect of increasing ctrl-G on ControlNet++ (Li et al., 2024a), a control adapter for StableDiffusion 1.5 (Rombach et al., 2022) that produces $\times512$ resolution text-prompt-conditioned images. Fig.˜14 illustrates qualitative improvements in control consistency (the shape of vegetation on the building, the curve of the window), whilst also quantitatively showing the control consistency vs FID tradeoff that ctrl-G introduces, aligning with earlier results in Fig.˜4. We use the evaluation code of Li et al. (2024a) [^8] https://github.com/liming-ai/ControlNet_Plus_Plus and calculate FID between the validation and generation images.
Figure 14: Ctrl-G can be applied to pre-trained opensource control adapters out of the box without the need for additional training. Left: Qualitative effect of increasing ctrl-G for ControlNet++ – consistency can be visually improved. Right: Quantitative tradeoff between consistency (segmentation mIoU) and quality (FID).